# Gaussian consensus processes and their Lyapunov exponents

### v1

@@ 3 Analysis of Model B
We claim that the coefficients $-\gamma\overline{Z}(t)=-\gamma\left(Z(t)-\mathbf{1}_{N,1}\mu^{Z}(t)\right)$ and $T(t)$ are globally Lipschitz functions of the state $Z(t)$. The existence and uniqueness of a strong solution follows from this claim by a well-known theorem of Itô (see for example [12, Thm IV.3.1] or [13, Thms 5.2.5 and 5.2.9]). Since all norms on a given finite-dimensional space of matrices are equivalent, it does not matter which norm we use to check the global Lipschitz condition. It is easiest to use the Frobenius norm. The Frobenius norm $\|A\|_{2}$ of an $m$ by $n$ (real or complex) matrix $A$ is simply the square root of the sum of the squares of the absolute values of its entries.
The absolute value of a rectangular real or complex matrix $A$ is the unique non-negative definite Hermitian square root of $A^{*}A$, where $A^{*}$ denotes the Hermitian conjugate of $A$. Note that $T(t)=\frac{1}{N}\left|\overline{Z}(t)\right|$. We apply the Araki-Yamagami inequality (originally proven in [2, Thm 1] for operators on a Hilbert space, but see [3, Thm VII.5.7, eqn VII.39] for a version that applies to arbitrary rectangular real or complex matrices). This says that the map that sends any matrix $A$ to its absolute value $|A|$ is Lipschitz with constant $\sqrt{2}$ with respect to the Frobenius norm:
@@ 3.1 Proof of Theorem 4
Let $W$ be an $N-1$ by $N-1$ matrix of independent standard Brownian motions and let $R$ be a continuous semimartingale adapted to the same filtration as $W$, such that the columns of $R$ are almost surely orthogonal. Let $S$ be the solution of the matrix Itô equation $dS=dWR$ with initial value $S(0)=\mathbf{0}_{N-1,d}$. Then $S$ is an $N-1$ by $d$ matrix of independent Brownian motions, also adapted to the same filtration as $W$.
@@
Let the processes $\hat{W}$, $H$, $\overline{Z}$, $Z$, $F$ and $T$ be as constructed in the statement of Theorem 4. We begin by showing that the process $VZ$, which takes its values in the space of $N-1$ by $d$ matrices, and is equal to $V\overline{Z}$, satisfies the following Stratonovich form SDE:
@@
where $W$ is an $N-1$ by $N-1$ matrix of independent standard Brownian motions. This follows on applying the Leibniz rule to the defining equation for $\overline{Z}$, which we recall:
@@
We see that the passage from Stratonovich to Itô form merely requires a modification of the coefficient of the drift of $VZ$ towards 00.
@@
$VZ(t)=V\overline{Z}(t)=e^{\gamma^{\prime}t}VV^{T}G(t/N)VZ(0)=e^{\gamma^{\prime}t}G(t/N)VZ(0)\,,$
@@
Since $G(t/N)\in GL(n,\mathbb{R})$ for all $t$, we find that $G(t/N)VZ(0)$ has the same rank as $VZ(0)$, namely $d$, and hence $\mathrm{Cov}^{Z}(t)$ is strictly positive definite, and so $T(t)$ is non-singular. Since the positive definite square-root map is analytic on the space of positive definite symmetric matrices, $T(t)$ is a continuous semimartingale, and so is $T(t)^{-1}$. Define
@@
which is to say that $R(t)$ is an $N-1$ by $d$ matrix with orthogonal columns. Moreover, $R$ is a continuous semimartingale adapted to the same filtration as $W$. Now we apply Lemma 14 to obtain an $N-1$ by $d$ matrix $S$ of independent Brownian motions such that
@@
$(d(VZ)(t))^{T}d(VZ)(t)=\frac{1}{\sqrt{N}}(d(VZ)(t))^{T}dW(t)VZ(t)=\frac{N-1}{N}(VZ(t))^{T}VZ(t)=(N-1)\mathrm{Cov}^{Z}(t)\,.$
@@
where $S(t)$ is some $N-1$ by $d$ matrix of independent Brownian motions adapted to the same filtration as $W$. Another calculation similar to the proof of Lemma 14 shows that because $R(t)$ has orthogonal columns, we may replace $dS(t)^{T}R(t)$ by a $d$ by $d$ matrix $U$ of independent standard Brownian motions adapted to the same filtration as $W$. We obtain
@@
With extra work we believe that it should be possible to show that the mapping from $Z$ to $\mathrm{Cov}^{Z}$ is a strong lumping, which is to say that for any time $s>0$, conditional on $\mathrm{Cov}^{Z}(s)$, the future process $\left(\mathrm{Cov}^{Z}(t)\right)_{t\geq s}$ and $Z(s)$ are independent.

### v2

@@ 3 Analysis of Model B
We claim that the coefficients $\gamma\overline{Z}(t)=-\gamma\left(Z(t)-\mathbf{1}_{N,1}\mu^{Z}(t)\right)$ and $T(t)$ are globally Lipschitz functions of the state $Z(t)$. The existence and uniqueness of a strong solution follows from this claim by a well-known theorem of Itô (see, for example, [12, Thm IV.3.1] or [13, Thms 5.2.5 and 5.2.9]). Since all norms on a given finite-dimensional space of matrices are equivalent, it does not matter which norm we use to check the global Lipschitz condition. It is easiest to use the Frobenius norm. The Frobenius norm $\|A\|_{2}$ of an $m$ by $n$ (real or complex) matrix $A$ is simply the square root of the sum of the squares of the absolute values of its entries.
The absolute value $|A|$ of a rectangular real or complex matrix $A$ is the unique non-negative definite Hermitian square root of $A^{*}A$, where $A^{*}$ denotes the Hermitian conjugate of $A$. Note that $T(t)=\frac{1}{N}\left|\overline{Z}(t)\right|$. We apply the Araki-Yamagami inequality (originally proven in [2, Thm 1] for operators on a Hilbert space, but see [3, Thm VII.5.7, eqn VII.39] for a version that applies to arbitrary rectangular real or complex matrices). This says that the map that sends any matrix $A$ to its absolute value $|A|$ is Lipschitz with constant $\sqrt{2}$ with respect to the Frobenius norm:
@@ 3.1 Proof of Theorem 4
In the proof of Theorem 4 we will apply the results of [23] taking n=N1n=N-1. We will also need the following simple result.
@@
Let $W$ be an N1N-1 by N1N-1 matrix of independent standard Brownian motions and let $R$ be a continuous semimartingale adapted to the same filtration as $W$, taking its values in N1N-1 by $d$ real matrices, such that the columns of $R$ are almost surely orthogonal. Let $S$ be the solution of the matrix Itô equation $dS=dWR$ with initial value $S(0)=\mathbf{0}_{N-1,d}$. Then $S$ is an N1N-1 by $d$ matrix of independent Brownian motions, adapted to the same filtration as $W$.
@@
Let the processes $\hat{W}$, $H$, $\overline{Z}$, $Z$, $F$ and $T$ be as constructed in the statement of Theorem 4. We begin by showing that the process $VZ$, which takes its values in the space of N1N-1 by $d$ matrices, and is equal to $V\overline{Z}$, satisfies the following Stratonovich form SDE:
@@
where $W$ is an N1N-1 by N1N-1 matrix of independent standard Brownian motions. This follows on applying the Leibniz rule to the defining equation for $\overline{Z}$, which we recall:
@@
We see that the passage from Stratonovich to Itô form merely requires a modification of the coefficient of the drift of $VZ$ towards 0.
@@
$VZ(t)=V\overline{Z}(t)=e^{-\gamma^{\prime}t}VV^{T}G(t/N)VZ(0)=e^{-\gamma^{\prime}t}G(t/N)VZ(0)\,$,
@@
Since $G(t/N)\in GL(N-1,\mathbb{R})$ for all $t$, we find that $G(t/N)VZ(0)$ has the same rank as $VZ(0)$, namely $d$, and hence $\mathrm{Cov}^{Z}(t)$ is strictly positive definite, and so $T(t)$ is non-singular. Since the positive definite square-root map is analytic on the space of strictly positive definite symmetric matrices, $T(t)$ is a continuous semimartingale, and similarly, since matrix inversion is analytic away from the singular matrices, $T(t)^{-1}$ is also a continuous semimartingale. Define
@@
which is to say that $R(t)$ is an N1N-1 by $d$ matrix with orthogonal columns. Moreover, $R$ is a continuous semimartingale adapted to the same filtration as $W$. Now we apply Lemma 12 to obtain an N1N-1 by $d$ matrix $S$ of independent Brownian motions such that
@@
$(d(VZ)(t))^{T}d(VZ)(t)$ $=\frac{1}{\sqrt{N}}(d(VZ)(t))^{T}dW(t)VZ(t)$
$=\frac{N-1}{N}(VZ(t))^{T}VZ(t)=(N-1)\mathrm{Cov}^{Z}(t)\,.$
@@
where $S(t)$ is some N1N-1 by $d$ matrix of independent Brownian motions adapted to the same filtration as $W$. Another calculation similar to the proof of Lemma 12 shows that because $R(t)$ has orthogonal columns, we may replace $dS(t)^{T}R(t)$ by a $d$ by $d$ matrix $U$ of independent standard Brownian motions adapted to the same filtration as $W$. We obtain
@@
With extra work we believe that it should be possible to show that the mapping from $Z$ to $\mathrm{Cov}^{Z}$ is a strong lumping, which is to say that for any time $s>0$, conditional on $\mathrm{Cov}^{Z}(s)$, the future process $\left(\mathrm{Cov}^{Z}(t)\right)_{t\geq s}$ and $Z(s)$ are independent. We expect this by analogy with Model A, which does have the corresponding strong lumping property, as we explained after the statement of Proposition 1.
4 Appendix
Recall the definition of the function $\phi_{m}$ in equation (3).
Lemma 13.
For all $m>0$ we have
Proof.
where $\xi\sim\mathrm{Poi}(x)$. Note also that
The expression for $\phi_{m}^{\prime}(x)$ can be simplified further:
Iterating this formula $\ell$ times, we obtain that
Recalling that $\phi_{m}(0)=\psi(m)$, we obtain
To get the second expression for $\phi_{m}(x)$, note that
and thus
$\int_{0}^{1}\frac{(1-e^{-xs})(1-s)^{m-1}}{s}\mathrm{d}s$ $=-\int_{0}^{1}\left(\sum_{k=1}^{\infty}\frac{(-x)^{k}s^{k-1}}{k!}\right)(1-s)^{m-1}\mathrm{d}s$
$=-\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\,\Gamma(k)}\int_{0}^{1}s^{k-1}(1-s)^{m-1}\mathrm{d}s$
$=-\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\,\Gamma(k)}\frac{\Gamma(k)\,\Gamma(m)}{\Gamma(m+k)}=-\sum_{k=1}^{\infty}\frac{(-x)^{k}}{k\cdot m(m+1)\dots(m+k-1)}.$
Proof of Proposition 3.
In case of an integer m=(N1)/2m=(N-1)/2, the expression obtained in (50) can be further simplified to
Let
Then
and
$=\sum_{j=1}^{m-1}\frac{(j-1)!}{(-x)^{j}}\sum_{i=j}^{m-1}\binom{i-1}{j-1}+\sum_{i=1}^{m-1}\frac{1}{i}+O(x)=\sum_{j=1}^{m-1}\frac{(j-1)!}{(-x)^{j}}\binom{m-1}{j}+\sum_{i=1}^{m-1}\frac{1}{i}+O(x)$
$=B+\sum_{i=1}^{m-1}\frac{1}{i}+O(x)$
$\lim_{x\downarrow 0}(A-B)=\sum_{i=1}^{m-1}\frac{1}{i},$
which together with Theorem 1 yields the statement of the claim. ∎
Remark 9.
Lemma 14.
If $m\geq 1$, then as $x\to\infty$
$\phi_{m}(x)=\log x+\frac{m-1}{x}+o\left(\frac{1}{x}\right).$
Proof.
From the integral representation
$\phi_{m}(x)-\psi(m)$ $=\int_{0}^{1}\frac{(1-e^{-xs})(1-s)^{m-1}}{s}\mathrm{d}s$
$=\int_{0}^{1}\frac{1-e^{-xs}}{s}\mathrm{d}s-(m-1)\int_{0}^{1}1-e^{-xs}\mathrm{d}s$
$+\int_{0}^{1}\frac{1-e^{-xs}}{s}\left[(1-s)^{m-1}-1+(m-1)s\right]\mathrm{d}s$
$=\gamma+\log x+\mathrm{Ei}_{1}(x)-(m-1)+\frac{(m-1)(1-e^{-x})}{x}+I$
where
$=\int_{0}^{1}\sum_{k=2}^{m-1}(-1)^{k}\binom{m-1}{k}s^{k-1}\mathrm{d}s-\int_{0}^{1}\frac{e^{-xs}}{s}\left[(1-s)^{m-1}-1+(m-1)s\right]\mathrm{d}s$
$=\sum_{k=2}^{m-1}\frac{(-1)^{k}}{k}\binom{m-1}{k}+R=(m-1)-\psi(m)-\gamma+R$
(from the Newton series for the digamma function) where
Note that $\sup_{s\in[0,1]}|g(s)|\leq m-1$ and by Peano’s formula
$g(s)=(m-1)(m-2)(1-\theta(s))^{m-3}\frac{s^{2}}{2!}.$
for some $0\leq\theta(s)\leq s$. Hence for $0\leq s\leq 1/2$ we have $|g(s)|\leq C_{m}s^{2}$ for some $C_{m}>0$. Consequently,
yielding
$\phi_{m}(x)-\log x$ $=\psi(m)+\gamma+\mathrm{Ei}_{1}(x)-(m-1)+\frac{m-1}{x}+I$
$=\mathrm{Ei}_{1}(x)+\frac{m-1}{x}+O\left(\frac{1}{x^{2}}\right)=\frac{m-1}{x}+o(1/x)$